A 1,340kg car moving at 20km/h is accelerated at a constant rate of 7m/s2 up to a speed of 88km/h. What is the total time required?
1
Expert's answer
2021-09-14T09:35:28-0400

The time it took to accelerate is


"t=\frac{\Delta v}{\Delta a}=\frac{(88-20)·1000/3600}{7}=2.7\text{ s}."